# Explainable Metrics for the Assessment of Neurodegenerative Diseases through Handwriting Analysis

### v2

@@ IV-B4 Writing Tasks Results
In this section, we present the results for the writing tasks’s specific metrics. As shown in Figure 16, AD participants take a much longer Average Time Between Writing than all other participants across all tasks, revealing a longer hesitation and intermitent pausing times, between words. CTL participants generally have lower variations in the pressure they use than ND participants, across all tasks, a potential indicator of more stability in steadiness in the writing. PD participants use less and less pressure as they write, in contrast to the AD and CTL participants, and they also use a larger area to write, significant only in the Copy Text Task, which correlates with previous findings about PD participants writing larger than others [1].
TABLE VIII: $p$-values and AUC of the specific metrics for all significant writing tasks. Significant p-values and AUC over 0.75 are bolded.
Writing Metrics AD vs CTL PD vs CTL PD vs PDM
@@
Copy change in pressure 0.810 0.508 0.043 0.624 0.115 0.573
Read press std 0.057 0.693 0.028 0.646 0.951 0.533
Text avg writing slope 0.028 0.524 0.060 0.599 0.710 0.504
avg time btwn write 0.009 0.710 0.264 0.583 0.665 0.702
press std 0.069 0.693 0.029 0.646 0.836 0.533
Copy area 0.112 0.612 0.044 0.542 0.897 0.562
Text avg writing slope 0.723 0.524 0.032 0.599 0.276 0.504
avg time btwn write 0.003 0.710 0.762 0.583 0.233 0.702
Freewrite press std 0.005 0.693 0.056 0.646 0.718 0.533
avg time btwn write 0.003 0.710 0.276 0.583 0.028 0.702
@@ IV-C Classification Results
This subsection discusses the classification results obtained from the different handwriting tasks analyzed in this study. The results are discussed by examining each group comparison individually. The models’ performance is evaluated regarding accuracy (ACC), F-1 score (F1), and AUC. Table IX presents results obtained using the Bagging classifier. The Appendix includes comprehensive experimental results from the RF and MLP classifiers.
TABLE IX: Classification results for pairwise group comparisons, using the Bagging model. The table reports experimental results in terms of accuracy (ACC), F1-score (F1), and AUC for each comparison. The best results for each metric for each group comparison are indicated in bold, while the second-best are italicized.
@@
When comparing AD against PD, the Copy Image and the Point Dominant tasks exhibit the best performance in terms of accuracy (91) and F1-Score (91). The Copy Image task involves visual perception and fine motor skills, the latter of which is particularly impaired in PD. Moreover, the Point Non-Dominant task reports the highest AUC score (94), indicating its efficacy in discriminating between the two conditions [45]. This is intriguing as mirror movement or mirroring dystonia have been described in Task Specific Dystonia and other movement disorders as a way to unmask and characterize latent dystonia features [46]; in our PD cohort, this task may be abnormal in PD subjects who are prone to the development of dystonic features - a finding generally though to be absent in AD.
@@
When contrasting AD with PDM, the highest accuracy (83) and the F1 score (83) are achieved in the Point Sustained task. Given the attentional deficits commonly observed in AD, this task’s effectiveness in discriminating between AD and CTL and PDM groups is plausible.
@@ V Discussion and Conclusion
This study presents a composite array of handwriting interpretable metrics and examines their efficacy in distinguishing between various NDs, including Alzheimer’s disease and Parkinson’s disease. To accomplish this, we introduce a novel dataset recorded from 113 participants at the Johns Hopkins Hospital. This dataset comprises 46 healthy controls and 73 individuals diagnosed with various neurodegenerative conditions. The dataset includes recordings of 14 handwriting tasks conducted on a tablet, complemented by synchronized speech recordings and eye-tracking data (which are not considered in this study), resulting in 1840 handwriting files. Our investigation encompasses diversified metrics, both task-agnostic and task-specific. We aim to assess their ability to effectively differentiate between different groups of participants across various NDs.
First, we extracted metrics agnostic of the task at hand, to measure dynamic properties of the handwriting, such as the hesitations and swiftness through the speed of writing, the noise of the signal through entropies, or the hesitations by time measures. Our findings demonstrate that task-agnostic metrics contribute significantly to the discrimination of NDs, both quantitatively and qualitatively. Quantitatively, a substantial portion of the derived task-agnostic metrics (53 out of 54) exhibit significant differences between at least two groups of participants. Qualitatively because multiple metrics are very discriminant between groups (p-values $\leq 10^{-7}$), consistent across tasks of the same category, and consistent with previous findings. However, mainly due to the limited sample of participants, not all expected handwriting signs associated to each of the ND at hand have been successfully observed through those task-agnostic metrics. For example, none of the metrics really grasped a significant micrographia [11] from PD participants.
Because of the obvious limits of the generalist approach, we also explored a task-specific approach, where metrics were designed to capture specific signs for each group of tasks.
For AD participants, the extracted metrics grasped higher variations for the Point tasks than the CTL participants, surely due to fatigue, a lower speed and a higher duration when drawing spirals, more variations of the pressure applied and more time between each word written when writing paragraphs, as well as when answering arythmetic questions.
For PD participants, we observe a sligthly higher radius than CTL participants for the point tasks, draw less loops of a spiral for the same duration than CTL, while being significantly slower than PDM participants when drawing spirals. When writing, both PD and PDM groups have more variations in pressure on all tasks than CTL participants.
However, the comparisons between tasks brought almost no significant difference between controls and ND participants, which will require further efforts to highlight.
We can still observe that while the PDM group is younger than the others, the extracted metrics behave between the PD and CTL groups most of the time, showing the proposed metrics exhibit a robustness to the age variations.
The use of the proposed set of metrics for binary classification yielded interesting results for all the groups of tasks. Both Drawing tasks and Point tasks are very discriminative for AD participants (up to 91% Accuracy against the PD group, up to 87% Accuracy against the CTL group). Meanwhile, the Spiral tasks are more effective in discriminating PDM participants, in particular the Spiral PaTaKa task that shows 73% accuracy against PD, 73% against AD and 77% against CTL. Across all tasks, AD discrimination against CTL is almost always more accurate than PD against CTL, which confirm our hypothesis that handwriting in general contains valuable informations for AD assessement and characterization. In general, the variations of performances within a task group are non-negligeable, which is supprising given that the differences between similar tasks explored in the section III-E2 did not show any significance, except for Copy Text vs Copy Read Text, which mean those differences should be studied further, using more generalist systems to tackle groups of tasks at a time.
This study presents the handwriting component of a newly developed multimodal and synchronized dataset designed for the digital characterization and assessment of NDs. A set of metrics, both generalizable and task-specific, were introduced to capture dynamic characteristics of handwriting across various tasks. The results demonstrate that several of these metrics effectively discriminate between different participant groups, maintain consistency across tasks with similar characteristics, and align with established indicators of their respective neurodegenerative conditions.
However, several limitations should be addressed. The sample size restricts generalization, especially considering the age difference between the PDM group and the others. The lack of significant differences between groups for some of the tasks indicates the need for more refined metrics. Although hand-crafted metrics were effective for some simple cases, keeping them explainable limits the complexity they can reach. In effect, not all tasks were able to produce discriminant metrics for all groups of participants, and not all expected behaviors usually observed in handwriting were observed through the proposed handcrafted metrics (for example, micrographia [11] was not observed in any of the tasks for PD participants). Future work should prioritize developing machine learning models that better capture the variations in the most complex patterns observed in handwriting.
Given the multimodal nature of the dataset presented, the integration of handwriting with synchronized speech and eye-tracking data offers a compelling avenue for enhanced assessment performance, which is out of the scope of this study. Future efforts will explore the potential benefits of integrating all available modalities to enhance performance on the task.
Real-world deployment will require portable, user-friendly tools. This work lays the foundation for handwriting analysis as a clinically impactful method for ND characterization, but presents limits that will be addressed in future works.

### v3

@@ IV-B4 Writing Tasks Results
In this section, we present the results for the writing tasks's specific metrics. As shown in Figure 14, AD participants take a much longer Average Time Between Writing than all other participants across all tasks, revealing a longer hesitation and intermitent pausing times, between words. CTL participants generally have lower variations in the pressure they use than ND participants, across all tasks, a potential indicator of more stability in steadiness in the writing. PD participants use less and less pressure as they write, in contrast to the AD and CTL participants, and they also use a larger area to write, significant only in the Copy Text Task, which correlates with previous findings about PD participants writing larger than others [1].
TABLE VII: $p$-values and AUC of the specific metrics for all significant writing tasks. Significant p-values and AUC over 0.75 are bolded.
Writing Metric AD vs CTL CTL vs PD PD vs PDM
@@
Copy Read Text press std 0.004 0.729 <<0.001 0.700 0.602 0.542
avg time btwn write 0.012 0.639 0.303 0.539 0.477 0.674
Copy Text press std 0.027 0.729 0.017 0.700 0.508 0.542
avg time btwn write 0.004 0.639 0.835 0.539 0.495 0.674
avg x dist btwn write 0.005 0.617 0.459 0.507 0.967 0.647
Freewrite press std <<0.001 0.729 <<0.001 0.700 0.628 0.542
avg time btwn write 0.019 0.639 0.539 0.539 0.046 0.674
@@ IV-C Classification Results
This subsection discusses the classification results obtained from the different handwriting tasks analyzed in this study. The results are discussed by examining each group comparison individually. The models' performance is evaluated in terms of accuracy (ACC), F-1 score (F1), and AUC. Table VIII presents results obtained using the Bagging classifier. Since the BG classifier consistently outperformed the other two classifiers we tested—RF and MLP—in terms of accuracy, F1-score, and AUROC in the different tasks considered, we present and discuss the results obtained with this classifier in the main text for the sake of simplicity. The experimental results for RF and MLP are provided in the Appendix.
TABLE VIII: Classification results for pairwise group comparisons, using the Bagging model. The table reports experimental results in terms of accuracy (ACC), F1-score (F1), and AUC for each comparison. The best results for each metric for each group comparison are indicated in bold, while the second-best are italicized.
@@
When comparing AD against PD, the Copy Image and the Point Dominant tasks exhibit the best performance in terms of accuracy (91) and F1-Score (91). The Copy Image task involves visual perception and fine motor skills, the latter of which is particularly impaired in PD. Moreover, the Point Non-Dominant task reports the highest AUC score (94), indicating its efficacy in discriminating between the two conditions [53]. This is intriguing as mirror movement or mirroring dystonia have been described in Task Specific Dystonia and other movement disorders as a way to unmask and characterize latent dystonia metrics [54]; in our PD cohort, this task may be abnormal in PD subjects who are prone to the development of dystonic metrics - a finding generally thought to be absent in AD.
@@
When contrasting AD with PDM, the highest accuracy (83) and the F1 score (83) are achieved in the Point Sustained task. Given the attentional deficits commonly observed in AD, this task's effectiveness in discriminating between AD and CTL and PDM groups is plausible.
@@
IV-D Correlation of the metrics with UPDRS-III and MoCa scores
To further explore the clinical relevance of the extracted handwriting metrics, we computed the Pearson correlation coefficient between each of the 54 task-agnostic metrics and the MoCA score [34], which evaluates global cognitive functioning, and the motor section of the Unified Parkinson's Disease Rating Scale (UPDRS-III) [35]. Only 13 out of 54 were significantly correlated with UPDRS-III scores ($p<0.05$), but none of them had $p$-values below 0.001, suggesting limited direct association between our handwriting metrics and motor symptom severity as assessed by UPDRS-III. However, a categorical examination of each subdomain of the UPDRS-III reveal a great variation of correlations depending on the domains, as show in Table IX. Asexpected, the tremor components are much more mesurable through handwriting than the postural and gait ones. The complete list of metrics with their individual significance is available in the appendix.
TABLE IX: Number of metrics over the 54 task-agnostic metrics produced being correlated with MoCa score, UPDRS-III score and its sub-components.
Score $p<0.05$ $p<0.01$ $p<0.001$ $p<0.0001$
MoCa [34] 33 22 15 9
UPDRS-III [35] 13 10 0 0
Rigidity 4 0 0 0
Upper bradykinesia 0 0 0 0
Lower bradykinesia 22 0 0 0
Arising from chair 2 0 0 0
Gait 20 4 0 0
Freeze of gait 8 2 2 2
Postural stability 0 0 0 0
Posture 1 1 0 0
Gait and posture composite 0 0 0 0
Kinetic tremor 19 5 0 0
Postural Tremor 15 6 0 0
Resting tremor 27 23 13 2
Global tremor 30 23 12 4
Non-tremor components 0 0 0 0
In contrast, a substantial number of metrics (33 out of 54) were significantly correlated with MoCA scores ($p<0.05$), indicating a strong relationship between handwriting behavior and cognitive performance. Among these, 9 metrics exhibited highly significant correlations with $p$-values below 0.0001, underscoring the potential of handwriting-based metrics as sensitive indicators of cognitive impairment.
@@ V Discussion and Conclusion
The primary goal of this study was to explore the utility of handwriting as a digital biomarker for the assessment and characterization of neurodegenerative diseases (NDs), including Alzheimer's disease (AD), Parkinson’s disease (PD), and related Parkinsonian disorders (PDM). We introduced a novel dataset collected at Johns Hopkins Hospital from 113 participants (46 healthy controls and 73 ND patients) who completed 14 structured handwriting tasks on a digital tablet. This rich dataset comprises 1840 handwriting recordings and is part of a broader multimodal acquisition effort that also includes synchronized speech and eye-tracking data.
We proposed and evaluated a large array of interpretable handwriting metrics designed to capture motor, cognitive, and behavioral markers from handwriting signals. These metrics were categorized as either task-agnostic (general signal descriptors applicable across tasks) or task-specific (metrics tailored to the expected behavior in particular task types). Our aim was to assess the ability of these metrics to distinguish between ND subtypes and healthy controls using statistical analysis and binary classification experiments, while ensuring clinical interpretability and relevance.
V-A Task-Agnostic and Task-Specific Metrics: Findings and Limitations
Our analysis of task-agnostic metrics, those computed independently of the specific task performed, demonstrated that a substantial number (53 out of 54) significantly differentiated at least one group comparison (e.g., PD vs. control, AD vs. PDM). These metrics included speed, acceleration, angular velocity, entropy, and timing-related metrics. Notably, these general metrics were consistent across task categories and aligned with known ND symptomatology. For example, reduced speed and increased entropy were particularly salient in AD participants, aligning with executive dysfunction and impaired motor planning.
However, the task-agnostic approach also revealed limitations. While general metrics captured high-level differences, they occasionally failed to detect hallmark disease-specific signs. Most notably, micrographia, a classic symptom of PD, was not captured in a statistically significant way. This suggests that some ND-specific metrics may only manifest under conditions that tightly control for task type, hand dominance, or cognitive load.
To address this limitation, we introduced task-specific metrics, tailored to the intent and structure of each task. These revealed more nuanced and disease-relevant patterns. AD participants showed longer in-air times, slower execution, and greater pressure variability, particularly during cognitively demanding tasks such as paragraph writing and arithmetic. PD participants drew fewer spiral loops, moved more slowly during motor tasks, and exhibited greater pressure variability, overlapping with patterns observed in the PDM group. While the PDM group was younger, its behavior often fell between the PD and control groups, suggesting that our metrics are relatively robust to age differences.
Despite these strengths, cross-task comparisons (such as between dominant and non-dominant hand usage or between reading and non-reading conditions) showed limited discriminative power, with the exception of Copy Text versus Copy Read Text. These results suggest the need for more advanced methods to capture inter-task behavioral adaptations.
V-B Classification Performance and Observed Limitations
In this study, we prioritized classification using interpretable metrics, incorporating task-specific and general metrics across handwriting tasks. Our primary goal was to develop a model that provides meaningful insights into the motor and cognitive impairments associated with neurodegenerative diseases, ensuring that the extracted metrics remain clinically interpretable. This approach allows for a better understanding of the underlying mechanisms driving classification decisions, which is crucial for real-world applications in clinical settings. However, deep learning-based approaches, such as those leveraging raw handwriting images or spectrogram representations of handwriting signals, offer a promising alternative. In a recent study [55], we explored non-interpretable methods by applying CNN and CNN-BLSTM models to spectrogram representations of handwriting signals for neurodegenerative disease classification. Our results demonstrated that classification performance varied across handwriting tasks and spectrogram channel combinations, with CNN consistently outperforming CNN-BLSTM. While the current study focused on interpretable metrics, future work could investigate hybrid approaches integrating deep learning techniques with interpretable models to enhance classification performance while maintaining clinical transparency. This would allow for the benefits of deep learning’s feature extraction capabilities while ensuring that the results remain interpretable and useful for clinical decision-making.
V-C Clinical Applications and Implications
Handwriting analysis has strong potential as a non-invasive, low-cost, and accessible digital biomarker. It captures both motor and cognitive processes and can be easily deployed using commercial tablets or styluses. Clinically, such tools could support early screening, aid differential diagnosis, and monitor disease progression or response to treatment. The metrics developed in this study are interpretable and aligned with known clinical metrics, making them suitable for real-world use.
While handwriting analysis shows significant promise as a digital biomarker for neurodegenerative diseases, the current equipment setup remains a notable limitation for widespread clinical adoption. The data in this study were collected using a high-resolution graphics tablet capable of capturing position, pressure, and in-air movements with high precision. However, this hardware is relatively expensive and requires controlled conditions and trained personnel for operation. These factors restrict its applicability outside research environments and limit deployment in routine clinical settings or home-based assessments.
V-D Discussion
While the proposed metrics and classification results offer compelling evidence for the utility of handwriting in distinguishing neurodegenerative conditions, there are several limitations that affect generalizability. First, the models developed in this study are trained on a dataset collected under tightly controlled conditions with a homogeneous set of equipment and task protocols. Their applicability to new populations, clinical settings, or hardware setups remains to be validated. This lack of generalization could be addressed through the use of pre-trained systems, such as neural networks which can learn more transferable representations, or through the use of more diverse handwriting datasets.
Additionally, while handwriting captures a rich combination of motor and cognitive activity, it is only one window into neurodegeneration. The broader dataset collected in this study also includes speech and eye-tracking data, which have yet to be analyzed in combination with handwriting. Multimodal approaches are likely to improve both sensitivity and specificity by modeling interactions across domains such as language, motor planning, and attention.
Finally, extending the current dataset by increasing the number of participants, balancing demographic variables, and adding more task repetitions will improve statistical power and enable more reliable model development and evaluation.
V-E General Conclusion and Future Directions
This study introduced a diverse set of interpretable metrics for handwriting analysis and demonstrated their value in differentiating between neurodegenerative conditions using data from structured tablet-based tasks. The results validate handwriting as a meaningful digital biomarker and provide a strong foundation for future research and clinical translation.
Building on this work, our future efforts will focus on four main directions. First, we plan to increase the size and diversity of our dataset to improve model robustness and enable the study of more subtle effects, such as early-stage impairments and cross-condition overlap. Second, we will integrate the available speech and eye-tracking data to explore multimodal models that can capture interaction effects across cognitive and motor channels. Third, we will investigate neural network-based solutions, including the use of pre-trained models, to leverage broader datasets and improve generalizability across populations and devices. Finally, we will evaluate the performance of our models on more portable and affordable equipment, with the goal of developing lightweight and scalable tools suitable for deployment in real-world clinical and home settings.
Together, these future directions aim to bridge the gap between research prototypes and practical, interpretable, and accessible tools for early detection and monitoring of neurodegenerative diseases.
Appendix A Additional Tables
TABLE X: Number of files and participants for each task and each experimental group. The symbol $\Sigma$ denotes the total count across experimental groups for a given task.
Task # Files # Participants
(Code) CTL PD PDM AD $\Sigma$ CTL PD PDM AD $\Sigma$
Copy Cube (CC) 38 25 14 39 116 36 22 12 20 90
Copy Image (CI) 42 37 16 42 137 39 34 14 21 108
Copy Image Memory (CM) 41 33 15 21 110 38 30 13 13 94
Copy Read Text (CR) 42 28 16 42 128 39 25 13 21 98
Copy Text (CT) 42 34 15 44 135 39 31 13 20 103
Draw Clock (DC) 43 27 15 41 126 40 24 13 20 97
Freewrite (F) 43 35 16 46 140 41 32 14 21 108
Numbers (N) 45 35 17 34 131 41 33 15 19 108
Point Dominant (PD) 44 36 16 43 139 41 32 13 21 107
Point Non-Dominant (PN) 43 36 17 43 139 40 33 14 21 108
Point Left (PL) 44 36 17 43 140 41 33 14 21 109
Point Right (PR) 43 36 16 43 138 40 32 13 21 106
Point Sustained (PS) 40 34 15 40 129 38 31 13 21 103
Spiral Dominant (SD) 41 36 16 45 138 38 33 14 21 106
Spiral Non-Dominant (SN) 41 36 15 45 137 40 33 13 21 107
Spiral Left (SL) 43 36 16 45 140 40 33 14 21 108
Spiral Right (SR) 39 36 15 45 135 38 33 13 21 105
Spiral Pataka (SP) 43 32 16 44 135 40 29 14 20 103
Total 588 464 219 569 1840 42 35 15 21 113
TABLE XI: Correlation of all the task-agnostic metrics with the MoCa score, UPDRS-III score and all its sub-components. 'n.s.' Is used for non significant features, * for p-values<0.05<0.05, ** for p-values<0.01<0.01, *** for p-values<0.001<0.001, and **** for p-values<0.0001<0.0001.
MoCa UPDRS Rigidity Upper Lower Arising Gait Freeze Postural Posture Gait and Kinetic Postural Resting Global Non
 [34] III brady brady from of gait stability posture tremor tremor tremor tremor tremor
 [35] kinesia kinesia chair composite components
in-air duration **** ** n.s. n.s. n.s. n.s. ** * n.s. ** n.s. n.s. n.s. * ** n.s.
in-air avg. X speed *** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * ** *** **** n.s.
in-air std X speed n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * **** **** n.s.
in-air amp X speed * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. *** *** n.s.
in-air avg. X acc. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** *** *** n.s.
in-air std X acc. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** *** **** n.s.
in-air amp X acc. ** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * * n.s.
in-air avg. Y speed *** n.s. n.s. n.s. n.s. * * n.s. n.s. n.s. n.s. * ** *** *** n.s.
in-air std Y speed n.s. n.s. n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. **** *** n.s.
in-air amp Y speed n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. *** *** n.s.
in-air avg. Y acc. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * ** *** *** n.s.
in-air std Y acc. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * *** **** n.s.
in-air amp Y acc. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. *** ** n.s.
in-air avg. speed ** n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** * *** *** n.s.
in-air std speed ** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** n.s. ** * n.s.
in-air amp speed *** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** n.s. ** ** n.s.
in-air avg. acc. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** * *** *** n.s.
in-air std acc. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s. ** * n.s.
in-air amp acc. ** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** n.s. ** ** n.s.
in-air std angular speed **** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s. * n.s.
in-air avg. angular speed **** n.s. n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. * * n.s. * n.s.
in-air std angular acc. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s. n.s. n.s.
in-air avg. angular acc. *** n.s. * n.s. n.s. n.s. ** n.s. n.s. n.s. n.s. * * n.s. n.s. n.s.
in-air shannon entropy **** n.s. n.s. n.s. n.s. * * * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet duration **** ** n.s. n.s. n.s. n.s. ** * n.s. n.s. n.s. n.s. n.s. * ** n.s.
on-tablet avg. X speed ** n.s. n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet std X speed * n.s. n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s.
on-tablet amp X speed n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet avg. X acc. n.s. n.s. n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s.
on-tablet std X acc. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet amp X acc. *** n.s. n.s. n.s. n.s. n.s. n.s. **** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet avg. Y speed * n.s. n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet std Y speed * n.s. n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s.
on-tablet amp Y speed n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet avg. Y acc. n.s. n.s. n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet std Y acc. n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet amp Y acc. ** n.s. n.s. n.s. * n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet avg. speed n.s. n.s. * n.s. * n.s. * n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s.
on-tablet std speed n.s. n.s. n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet amp speed n.s. * n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet avg. acc. n.s. n.s. n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s.
on-tablet std acc. n.s. * n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s.
on-tablet amp acc. n.s. * n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet std angular speed **** n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s.
on-tablet avg. angular speed **** ** * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet std angular acc. **** ** n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. * ** ** n.s.
on-tablet avg. angular acc. *** ** n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. n.s. ** n.s. * n.s.
on-tablet avg. dP/dt ** n.s. n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. n.s. n.s. * n.s. n.s.
on-tablet std dP/dt * ** n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. * n.s. ** ** n.s.
on-tablet amp dP/dt * ** n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** ** n.s.
on-tablet avg. d2P/dt2 * ** n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** ** n.s.
on-tablet std d2P/dt2 * ** n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. * n.s. ** ** n.s.
on-tablet amp d2P/dt2 * ** n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** ** n.s.
on-tablet shannon entropy **** n.s. n.s. n.s. n.s. n.s. ** **** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
TABLE XII:
Task AD vs PD PD vs PDM AD vs CTL AD vs PDM PD vs CTL PDM vs CTL
ACC F1 AUC ACC F1 AUC ACC F1 AUC ACC F1 AUC ACC F1 AUC ACC F1 AUC
Random Forest
Copy Cube 64 63 73 53 52 58 63 64 70 67 65 69 64 63 66 65 62 55
Copy Image 83 83 89 53 46 62 79 79 84 78 78 80 52 52 49 58 55 64
Copy Image Memory 77 77 82 57 55 56 62 62 60 64 64 65 69 69 75 62 59 50
Copy Read Text 72 71 78 57 56 54 63 63 70 79 78 82 65 63 66 67 64 79
Copy Text 82 82 90 77 76 77 79 79 85 77 75 81 61 61 61 64 60 64
Draw Clock 78 77 88 50 49 52 77 78 88 79 78 68 60 60 56 64 64 67
Freewrite 73 73 73 59 56 70 71 71 77 79 79 78 59 58 55 75 74 79
Numbers 66 65 70 57 49 46 68 68 73 51 47 48 50 50 52 57 50 55
Point Dominant 88 88 93 43 41 39 83 82 84 82 82 80 62 62 62 59 57 53
Point Non-Dominant 88 88 95 65 64 56 84 83 89 78 76 73 68 68 71 67 66 67
Point Sustained 89 89 93 63 59 51 81 81 90 83 82 79 65 65 66 62 62 57
Spiral Dominant 72 71 71 60 55 56 68 68 68 73 70 58 49 48 56 69 67 64
Spiral Non-Dominant 63 62 63 68 67 68 70 69 63 75 74 67 45 43 41 64 62 59
Spiral PaTaKa 68 66 69 66 65 46 58 57 52 73 71 71 60 59 62 79 78 69
Multi Layer Perceptron
Copy Cube 78 78 81 53 51 52 56 56 57 69 69 69 57 57 59 52 53 51
Copy Image 84 84 90 67 63 56 72 72 82 69 68 67 59 59 62 49 48 48
Copy Image Memory 65 64 76 57 56 59 55 55 51 56 56 66 57 57 64 56 57 51
Copy Read Text 65 63 66 32 33 33 66 66 73 78 77 87 61 59 60 69 66 79
Copy Text 77 78 84 65 63 69 73 73 78 86 85 92 58 58 62 71 71 66
Draw Clock 73 72 79 59 59 51 79 79 88 77 76 69 54 54 61 52 53 50
Freewrite 60 59 63 59 56 50 62 62 72 75 75 77 45 45 44 73 72 75
Numbers 46 46 44 39 38 40 56 56 70 45 45 62 41 41 46 57 55 53
Point Dominant 88 88 92 57 53 57 81 81 84 74 71 75 62 62 73 64 62 65
Point Non-Dominant 83 83 89 57 55 57 84 83 82 75 72 72 59 58 68 64 64 65
Point Sustained 88 88 91 65 63 67 87 87 85 81 81 80 65 65 68 65 64 59
Spiral Dominant 67 66 74 47 42 38 59 58 58 78 75 51 56 56 55 74 74 64
Spiral Non-Dominant 61 59 60 75 73 59 69 68 69 73 66 70 48 48 49 57 51 52
Spiral PaTaKa 66 66 67 70 70 54 50 50 52 79 77 71 58 58 52 77 76 71